# Improved Margin Generalization Bounds for Voting Classifiers

### v1

@@ 3 Upper bound
In this section we prove Lemma 4, which is used in a union bound over suitable $\gamma_{0},\gamma_{1}$ and $\tau_{0}$ and $\tau_{1},$ giving the margin bound in Theorem 1 for all $f\in\operatorname{\Delta({\mathcal{H}})}$ and margins $0<\gamma\leq 1$
@@
Now, for $a>0,$ $x-\sqrt{ax}$ is increasing for $x\geq a/4,$ since it has derivative $1-\frac{a}{2\sqrt{ax}}$. Thus, for $\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\geq\tau_{1}+64\beta=\tau_{1}+64\left(\sqrt{\frac{\tau_{1}\cdot 2\ln{\left(\frac{e}{\delta}\right)}}{m}}+\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}\right)\geq\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}$, we have by the above with $a=\frac{2\ln{\left(e/\delta\right)}}{m},$ that with probability at least $1-\frac{\delta}{e}$ over ${\mathbf{S}}^{\prime},$ we have
@@
$\geq\tau_{1}+32\left(\sqrt{\frac{\tau_{1}\cdot 2\ln{\left(\frac{e}{\delta}\right)}}{m}}+\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}\right)=\tau+32\beta,$
@@
By ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$ being i.i.d., we may see ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$ as being drawn in the following way: First, we draw $\bar{{\mathbf{S}}}\sim{\mathcal{D}}^{2m}$ and then let ${\mathbf{S}}$ be created by drawing $m$ times without replacement from $\bar{{\mathbf{S}}}$ and ${\mathbf{S}}^{\prime}=\bar{{\mathbf{S}}}\backslash{\mathbf{S}}$. We denote this way of drawing ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$ from $\bar{{\mathbf{S}}}$ as ${\mathbf{S}},{\mathbf{S}}^{\prime}\sim\bar{{\mathbf{S}}}$. We then have that,
@@
$\leq\operatorname*{\mathbb{P}}_{\bar{{\mathbf{S}}}\sim{\mathcal{D}}^{2m}}\Bigg{[}\operatorname*{\mathbb{P}}_{{\mathbf{S}},{\mathbf{S}}^{\prime}\sim\bar{{\mathbf{S}}}}\Bigg{[}\exists\gamma\in\left[\gamma_{0},\gamma_{1}\right],\exists f\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}\right\rceil}:\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}\left(f\right)\leq\tau_{1},\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}(f)\geq\tau_{1}+32\beta\Bigg{]}\Bigg{]}$
$\leq\sup_{Z\in({\mathcal{X}}\times\{-1,1\})^{2m}}\operatorname*{\mathbb{P}}_{{\mathbf{S}},{\mathbf{S}}^{\prime}\sim Z}\bigg{[}\exists\gamma\in\left[\gamma_{0},\gamma_{1}\right],\exists f\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}\right\rceil}:\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}\left(f\right)\leq\tau_{1},\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}(f)\geq\tau_{1}+32\beta\bigg{]}.$
@@
Recall that $\beta=\left(\sqrt{\frac{\tau_{1}\cdot 2\ln{\left(\frac{e}{\delta}\right)}}{m}}+\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}\right).$ Let $2\mu$ denote the fraction of points in $Z$ where $f$ has less than $\gamma_{0}/2$-margin, i.e.,
@@
If $\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\frac{\gamma_{0}}{2}}(f)\leq\tau_{1}$, then by $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$, we have that
@@
Now if $\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}^{\frac{\gamma_{0}}{2}}(f)\geq\tau_{1}+32\beta=\geq\tau_{1}+32\Big{(}\sqrt{\frac{\tau_{1}\cdot 2\ln{\left(\frac{e}{\delta}\right)}}{m}}+\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}\Big{)}$ and using that we argued earlier that $x-\sqrt{ax}$ is increasing for $x\geq\frac{a}{4}$, we get that
@@
Now upper bounding the second term using by $a+\sqrt{ab}+b\leq 2(a+b)$ and $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b},$
@@ 4 Bound on infinity cover
In this section, we prove Lemma 5, where from Corollary 6 follows and gives us the covering bound needed as described in the proof sketch. To this end recall that we for a function class ${\mathcal{F}}$ and a point set $X=\{x_{1},\ldots,x_{m}\}\subset{\mathcal{X}}$ of size $m$ and precision $\gamma$ define ${\mathcal{N}}_{\infty}(X,{\mathcal{F}},\gamma)$ to be the minimal size of a $\gamma$-cover in infinity norm of ${\mathcal{F}}$ on $X,$ i.e. the smallest size of a set of functions $N$ satisfying $\min_{f^{\prime}\in N}\max_{x\in X}|f(x)-f^{\prime}(x)|\leq\gamma.$ Furthermore, we will use $\operatorname{\mathrm{Ln}}(x)=\ln(\max(x,e))$ for the truncated natural logarithm. We now state Lemma 5 and Corollary 6, and show how Lemma 5 implies Corollary 6.
@@
$\ln({\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil},\alpha\gamma))\leq\frac{c^{\prime}d}{\alpha^{2}\gamma^{2}}\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})})\operatorname{\mathrm{Ln}}{\left(\frac{8\alpha m\gamma^{2}}{cd}\right)}=\frac{c^{\prime}d}{\alpha^{2}\gamma^{2}}\operatorname{\Gamma}\left(\frac{8\alpha m\gamma^{2}}{cd}\right)$ (32)
@@
Let $c^{\prime},c$ be the constants from Lemma 5, and furthermore consider $\gamma^{2}\geq\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})})}{\alpha^{2}m}.$ To the end of invoking Lemma 5, notice that for $\varepsilon=1/(2\operatorname{\mathrm{Ln}}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})}))\leq 1/2,$ we have that $\frac{cd}{\alpha^{2}\varepsilon^{2}m}\leq\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})})}{\alpha^{2}m}\leq\gamma^{2},$ and we thus may invoke Lemma 5 with this $\varepsilon.$ We notice that for this $\varepsilon$ we have that
@@
where we in the inequality used the value of $\varepsilon$ and that $\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\varepsilon^{2}\gamma^{2}}{cd})}\leq\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})}.$ Thus we conclude that
@@
In what follows let, $\check{C}\geq 1$ and $\check{c}>0$ be the universal constants from Lemma 7, and $C\geq 1$ be the universal constant from Lemma 8. Furthermore, let $c$ denote the universal constant $c=\frac{16C}{\check{c}^{2}}$ and $c^{\prime}=\frac{16\check{C}C}{\check{c}^{2}}$ in Lemma 5. Since $\check{C}\geq 1$ we have $c^{\prime}\geq c.$
By Lemma 8, we have that $d^{\prime}=\operatorname{\mathrm{fat}}_{\check{c}\alpha\varepsilon\gamma}(\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil})$ is upper bounded by $\frac{4Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}$. Thus, we have that $d^{\prime}=\operatorname{\mathrm{fat}}_{\check{c}\alpha\varepsilon\gamma}(\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil})\leq m$, by the assumption $\gamma^{2}\geq\frac{cd}{\alpha^{2}\varepsilon^{2}m}$ and $c=\frac{16C}{\check{c}^{2}},$ and we may invoke Lemma 7 to obtain
@@
Next, we analyze the term $d^{\prime}\ln^{1+\varepsilon}\left(\frac{2m}{d^{\prime}\alpha}\right)$. We now notice that the function $x\ln^{1+\varepsilon}{\left(\frac{a}{x}\right)}$ has derivative $\ln^{\varepsilon}{\left(\frac{a}{x}\right)}\left(\ln{\left(\frac{a}{x}\right)}-1-\varepsilon\right)$, thus the derivative is positive for $\ln{\left(\frac{a}{x}\right)}\geq 1+\varepsilon$ or equivalently $\frac{a}{\exp{\left(1+\varepsilon\right)}}\geq x$, whereby we conclude that $x\ln^{1+\varepsilon}{\left(\frac{a}{x}\right)}$ is increasing for $\frac{a}{\exp{\left(1+\varepsilon\right)}}\geq x$. Now using this with $a=\frac{2m}{\alpha}\geq\frac{64Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}$, where we used that $\gamma^{2}\geq\frac{cd}{\alpha^{2}\varepsilon^{2}m}$, $c=\frac{16C}{\check{c}^{2}}$ and $\alpha<\frac{1}{2}$, and having the upper bound on $d^{\prime}=\operatorname{\mathrm{fat}}_{\check{c}\alpha\varepsilon\gamma}(\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil})$ of $\frac{4Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}\leq\frac{\frac{32Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}}{\exp{\left(1+\varepsilon\right)}}$, since $\frac{32}{\exp{\left(1+\varepsilon\right)}}\geq 4$, we conclude that $d^{\prime}=\operatorname{\mathrm{fat}}_{\check{c}\alpha\varepsilon\gamma}(\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil})\leq\frac{4Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}\leq\frac{a}{\exp{\left(1+\varepsilon\right)}}.$ This then implies by the above argued monotonicity of $x\ln^{1+\varepsilon}{\left(\frac{a}{x}\right)},$ with $a=\frac{2m}{\alpha}$ that
$\check{C}d^{\prime}\ln^{1+\varepsilon}{\left(\frac{2m}{d^{\prime}\alpha}\right)}\leq\check{C}\frac{4Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}\ln^{1+\varepsilon}{\left(\frac{2m}{\alpha}\frac{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}{4Cd}\right)}\leq\frac{c^{\prime}d}{\alpha^{2}\varepsilon^{2}\gamma^{2}}\ln^{1+\varepsilon}{\left(\frac{8\alpha m\varepsilon^{2}\gamma^{2}}{cd}\right)},$
where the last equivalently follows from $c=\frac{16C}{\check{c}^{2}}$ and $c^{\prime}=\frac{16\check{C}C}{\check{c}^{2}}$ (we notice that if $c^{\prime}$ is not at least $1$, we could set it equal to $1$ and still have an upper bound) which concludes the proof. ∎
@@
i.e. $f$ always being below $f_{\left\lceil\gamma\right\rceil}$ if $f_{\left\lceil\gamma\right\rceil}$ is less strictly less than $\gamma$ and $f$ always being above $f_{\left\lceil\gamma\right\rceil}$ if $f_{\left\lceil\gamma\right\rceil}$ is strictly larger than $-\gamma$. This implies that the function $f\in\operatorname{\Delta({\mathcal{H}})}$ generating $f_{\left\lceil\gamma\right\rceil}$ also satisfies:
@@
Furthermore, we note that for any $f\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil}$ and $f^{\prime}\in N,$ where $N$ is a minimal (in terms of size) $\alpha$-cover of $\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil}$ on $X$ in infinity norm, such that $f^{\prime}$ is $\alpha$-close to $\frac{f}{\gamma}\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil}/\gamma$, it holds that

### v2

@@ 3 Upper bound
In this section we prove Lemma 4, which is used in a union bound over suitable $\gamma_{0},\gamma_{1}$ and $\tau_{0}$ and $\tau_{1},$ giving the margin bound in Theorem 1 for all $f\in\operatorname{\Delta({\mathcal{H}})}$ and margins $0<\gamma\leq 1$ We now state [lem:marginbound].
@@
Now, for $a>0,$ $x-\sqrt{ax}$ is increasing for $x\geq a/4,$ since it has derivative $1-\frac{a}{2\sqrt{ax}}$. Thus, for $\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\geq\tau_{1}+64\beta=\tau_{1}+64\left(\sqrt{\frac{\tau_{1}\cdot 2\ln{\left(\frac{e}{\delta}\right)}}{m}}+\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}\right)\geq\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}$, we have by the above with $a=\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m},$ that with probability at least $1-\frac{\delta}{e}$ over ${\mathbf{S}}^{\prime},$ we have
@@
$\geq\tau_{1}+32\left(\sqrt{\frac{\tau_{1}\cdot 2\ln{\left(\frac{e}{\delta}\right)}}{m}}+\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}\right)=\tau_{1}+32\beta,$
@@
By ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$ being i.i.d., we may see ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$ as being drawn in the following way: First, we draw $\bar{{\mathbf{S}}}\sim{\mathcal{D}}^{2m}$ and then let ${\mathbf{S}}$ be created by drawing $m$ times without replacement from $\bar{{\mathbf{S}}}$ and ${\mathbf{S}}^{\prime}=\bar{{\mathbf{S}}}\backslash{\mathbf{S}}$. We denote this way of drawing ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$ from $\bar{{\mathbf{S}}}$ as $({\mathbf{S}},{\mathbf{S}}^{\prime})\sim\bar{{\mathbf{S}}}$. We then have that,
@@
$=\operatorname*{\mathbb{P}}_{\bar{{\mathbf{S}}}\sim{\mathcal{D}}^{2m}}\Bigg{[}\operatorname*{\mathbb{P}}_{({\mathbf{S}},{\mathbf{S}}^{\prime})\sim\bar{{\mathbf{S}}}}\Bigg{[}\exists\gamma\in\left[\gamma_{0},\gamma_{1}\right],\exists f\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}\right\rceil}:\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}\left(f\right)\leq\tau_{1},\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}(f)\geq\tau_{1}+32\beta\Bigg{]}\Bigg{]}$
$\leq\sup_{Z\in({\mathcal{X}}\times\{-1,1\})^{2m}}\operatorname*{\mathbb{P}}_{({\mathbf{S}},{\mathbf{S}}^{\prime})\sim Z}\bigg{[}\exists\gamma\in\left[\gamma_{0},\gamma_{1}\right],\exists f\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil 2\gamma_{1}\right\rceil}:\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma}\left(f\right)\leq\tau_{1},\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}(f)\geq\tau_{1}+32\beta\bigg{]}.$
@@
Recall that $\beta=\left(\sqrt{\frac{\tau_{1}\cdot 2\ln{\left(\frac{e}{\delta}\right)}}{m}}+\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}\right).$ Let $2\mu$ denote 2 times the fraction of points in $Z$ where $f$ has less than $\gamma_{0}/2$-margin, i.e.,
@@
If $\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\frac{\gamma_{0}}{2}}(f)\leq\tau_{1}$, we have that
@@
Now if $\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}^{\frac{\gamma_{0}}{2}}(f)\geq\tau_{1}+32\beta=\tau_{1}+32\Big{(}\sqrt{\frac{\tau_{1}\cdot 2\ln{\left(\frac{e}{\delta}\right)}}{m}}+\frac{2\ln{\left(\frac{e}{\delta}\right)}}{m}\Big{)}$ and using that we argued earlier that $x-\sqrt{ax}$ is increasing for $x\geq\frac{a}{4}$, we get that
@@
Now upper bounding the second term by using that $a+\sqrt{ab}+b\leq 2(a+b)$ and $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b},$
@@ 4 Bound on infinity cover
In this section, we prove the necessary result on the infinity cover of $\Delta({\mathcal{H}})_{\left\lfloor\gamma\right\rfloor}$. To this end we recall that we for a function class ${\mathcal{F}}$ and a point set $X=\{x_{1},\ldots,x_{m}\}\subset{\mathcal{X}}$ of size $m$ and precision $\gamma$ define ${\mathcal{N}}_{\infty}(X,{\mathcal{F}},\gamma)$ to be the minimal size of a $\gamma$-cover in infinity norm of ${\mathcal{F}}$ on $X,$ i.e. the smallest size of a set of functions $N$ satisfying $\min_{f^{\prime}\in N}\max_{x\in X}|f(x)-f^{\prime}(x)|\leq\gamma.$ Furthermore, we will use $\operatorname{\mathrm{Ln}}(x)=\ln(\max(x,e))$ for the truncated natural logarithm. We now state Lemma 5 and Corollary 6, and show how Lemma 5 implies Corollary 6.
@@
$\ln({\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil},\alpha\gamma))\leq\frac{c^{\prime}d}{\alpha^{2}\gamma^{2}}\operatorname{\mathrm{Ln}}^{2}\left(\operatorname{\mathrm{Ln}}{\left(\frac{8\alpha m\gamma^{2}}{cd}\right)}\right)\operatorname{\mathrm{Ln}}{\left(\frac{8\alpha m\gamma^{2}}{cd}\right)}=\frac{c^{\prime}d}{\alpha^{2}\gamma^{2}}\operatorname{\Gamma}\left(\frac{8\alpha m\gamma^{2}}{cd}\right)$ (32)
@@
Let $c^{\prime},c$ be the constants from Lemma 5, and furthermore consider $\gamma^{2}\geq\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})})}{\alpha^{2}m}.$ To the end of invoking Lemma 5, notice that for $0<\varepsilon=1/(2\operatorname{\mathrm{Ln}}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})}))\leq 1/2,$ we have that $\frac{cd}{\alpha^{2}\varepsilon^{2}m}=\frac{4cd\operatorname{\mathrm{Ln}}^{2}(\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})})}{\alpha^{2}m}\leq\gamma^{2},$ and we thus may invoke Lemma 5 with this $\varepsilon.$ We notice that for this $\varepsilon$ we have that
@@
where we in the inequality used the value of $\varepsilon$ and that $\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\varepsilon^{2}\gamma^{2}}{cd})}\leq\operatorname{\mathrm{Ln}}{(\frac{8\alpha m\gamma^{2}}{cd})}$ and $\ln$ being an increasing function. Thus we conclude that
@@
In what follows let, $\check{C}\geq 1$ and $\check{c}>0$ be the universal constants from Lemma 7, and $C\geq 1$ be the universal constant from Lemma 8. Furthermore, let $c$ denote the universal constant $c=\frac{16C}{\check{c}^{2}}$ and $c^{\prime}=\max(1,\frac{16\check{C}C}{\check{c}^{2}})$ in Lemma 5. Since $\check{C}\geq 1$ we have $c^{\prime}\geq c.$
By Lemma 8, we have that $d^{\prime}=\operatorname{\mathrm{fat}}_{\check{c}\alpha\varepsilon\gamma}(\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil})$ is upper bounded by $\frac{Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}$. Thus, we have that $d^{\prime}=\operatorname{\mathrm{fat}}_{\check{c}\alpha\varepsilon\gamma}(\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil})\leq m$, by the assumption $\gamma^{2}\geq\frac{cd}{\alpha^{2}\varepsilon^{2}m}$ and $c=\frac{16C}{\check{c}^{2}},$ and we may invoke Lemma 7 to obtain
@@
Next, we analyze the term $d^{\prime}\ln^{1+\varepsilon}\left(\frac{2m}{d^{\prime}\alpha}\right)$. We now notice that the function $x\ln^{1+\varepsilon}{\left(\frac{a}{x}\right)}$ for has derivative $\ln^{\varepsilon}{\left(\frac{a}{x}\right)}\left(\ln{\left(\frac{a}{x}\right)}-1-\varepsilon\right)$, thus the derivative is positive for $\ln{\left(\frac{a}{x}\right)}\geq 1+\varepsilon$ or equivalently $\frac{a}{\exp{\left(1+\varepsilon\right)}}\geq x$, whereby we conclude that $x\ln^{1+\varepsilon}{\left(\frac{a}{x}\right)}$ is increasing for $\frac{a}{\exp{\left(1+\varepsilon\right)}}\geq x$. Now using this with $a=\frac{2m}{\alpha}\geq\frac{64Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}$, where we used that $\gamma^{2}\geq\frac{cd}{\alpha^{2}\varepsilon^{2}m}$, $c=\frac{16C}{\check{c}^{2}}$ and $\alpha<\frac{1}{2}$, and having the upper bound on $d^{\prime}=\operatorname{\mathrm{fat}}_{\check{c}\alpha\varepsilon\gamma}(\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil})$ of $\frac{Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}\leq\frac{\frac{32Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}}{\exp{\left(1+\varepsilon\right)}}$, since $\frac{32}{\exp{\left(1+\varepsilon\right)}}\geq 1$, we conclude that $d^{\prime}=\operatorname{\mathrm{fat}}_{\check{c}\alpha\varepsilon\gamma}(\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil})\leq\frac{Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}\leq\frac{a}{\exp{\left(1+\varepsilon\right)}}.$ This then implies by the above argued monotonicity of $x\ln^{1+\varepsilon}{\left(\frac{a}{x}\right)},$ with $a=\frac{2m}{\alpha}$ that
$\check{C}d^{\prime}\ln^{1+\varepsilon}{\left(\frac{2m}{d^{\prime}\alpha}\right)}\leq\check{C}\frac{Cd}{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}\ln^{1+\varepsilon}{\left(\frac{2m}{\alpha}\frac{\check{c}^{2}\alpha^{2}\varepsilon^{2}\gamma^{2}}{Cd}\right)}\leq\frac{c^{\prime}d}{\alpha^{2}\varepsilon^{2}\gamma^{2}}\ln^{1+\varepsilon}{\left(\frac{8\alpha m\varepsilon^{2}\gamma^{2}}{cd}\right)},$
where the last equivalently follows from $c=\frac{16C}{\check{c}^{2}}$ and $c^{\prime}=\max(1,\frac{16\check{C}C}{\check{c}^{2}})$ (we notice that if $c^{\prime}$ is not at least $1$, we could set it equal to $1$ and still have an upper bound) which concludes the proof. ∎
@@
i.e. $f$ always being below $f_{\left\lceil\gamma\right\rceil}$ if $f$ is less strictly less than $\gamma$ and $f$ always being above $f_{\left\lceil\gamma\right\rceil}$ if $f$ is strictly larger than $-\gamma$. This implies that the function $f\in\operatorname{\Delta({\mathcal{H}})}$ generating $f_{\left\lceil\gamma\right\rceil}$ also satisfies:
@@
Furthermore, we note that for any $f\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil}$ and $f^{\prime}\in N,$ where $N$ is a minimal (in terms of size) $\alpha$-cover of $\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil}/\gamma$ on $X$ in infinity norm, such that $f^{\prime}$ is $\alpha$-close to $\frac{f}{\gamma}\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma\right\rceil}/\gamma$, it holds that